# Improved Constructions of Reed-Solomon Codes with Optimal Repair Bandwidth

### v1

@@ II-C Reed-Solomon codes as MSR codes
Let $E$ be an extension field of finite field $F$ with $E:F]=\ell$. Assume $\{\zeta_{1},\dots,\zeta_{\ell}\}$ and $\{\mu_{1},\dots,\mu_{\ell}\}$ are dual basis of $E$ over $F$, i.e., ${\rm tr}(\zeta_{i}\mu_{j})=\delta_{ij}$, where ${\rm tr}(\cdot)$ denote the trace function from $E$ to $F$ and $\delta_{ij}$ is the Kronecker delta. Then for any $\gamma\in E$,
@@
For a vector space $V$ over a field $F$ and a set of vectors $A=\left(a_{1},\ldots,a_{l}\right)\subset V$, let $\operatorname{Span}_{F}(A)=\left\{\sum_{i=1}^{l}\gamma_{i}a_{i},\gamma_{i}\in F\right\}$ be the span of $A$ over $F$.
@@
    Let $h(x)=\prod_{j\in[n]\setminus(R_{i}\cup\{i\})}(x-\alpha_{j})$, then for $0\leq t\leq s-1$, $(v_{1}\alpha_{1}^{t}h(\alpha_{1}),\dots,v_{n}\alpha_{n}^{t}h(\alpha_{n}))\in C^{\perp},$ hence $\sum_{j=1}^{n}c_{j}v_{j}\alpha_{j}^{t}h(\alpha_{j})=0$
@@
    Assume that $R_{i}=\{i_{1},\dots,i_{d}\}$, then
    $\sum_{j=1}^{d}c_{i_{j}}v_{i_{j}}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}})=0.$
@@
    can be used to recover $c_{i}$, as $\left\{e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i}))\right\}_{0\leq t\leq s-1,1\leq m\leq p_{i}}$ forms a basis of $\mathbb{E}$ over $F_{i}$.
@@
    Hence, from each helper node, for example, the $i_{1}$-th node in $R_{i}$, one only need to calculate $\{{\rm tr}_{i}(c_{i_{1}}v_{i_{1}}e_{m})\}_{1\leq m\leq p_{i}}$, i.e. $p_{i}$ symbols form $F_{i}$. Consequently, the total repair bandwidth is $dp_{i}=\frac{sp_{i}d}{d+1-k}$, meet the cut-set bound (1).
@@ III Main results
In this section, we generalize Lemma 2 by constructing an explicit $S$ satisfying (3) without assuming $p_{i}\equiv 1\pmod{s}$, thereby obtaining an RS–MSR code with smaller subpacketization $\ell=s\prod_{i=1}^{n}p_{i}$ for any choice of distinct primes $p_{1},\ldots,p_{n}>s$.
Lemma 3.
@@
Figure 1: Original ($p\times s$)
@@
Figure 2: Target ($s\times p$)
@@
where $b_{m}=\gcd(p,s)$.
Given a $p\times s$ array $A$, we partition its entries into square blocks by partitioning its index set
$\mathcal{I}(p,s)\triangleq[0,p-1]\times[0,s-1]$
into disjoint squares via the following deterministic recursive procedure.
@@
Let $h,w$ be positive integers and let
$\mathcal{I}(h,w)\triangleq[0,h-1]\times[0,w-1].$
We define a partition of $\mathcal{I}(h,w)$ into disjoint squares via the following recursive procedure. Let $t:=\min\{h,w\}$. Tile $\mathcal{I}(h,w)$ by placing as many $t\times t$ squares as possible along the longer side (left-to-right if $w\geq h$, and top-to-bottom if $h>w$). This leaves a (possibly empty) residual rectangle, denoted by $\mathcal{I}(h^{\prime},w^{\prime})$; if it is nonempty, apply the same procedure recursively to $\mathcal{I}(h^{\prime},w^{\prime})$. The resulting collection of squares forms a disjoint partition of $\mathcal{I}(h,w)$, called the Euclidean square partition of $\mathcal{I}(h,w)$ and denoted by $\mathcal{P}(h,w)$. Each square in $\mathcal{P}(h,w)$ is called a Euclidean square.
This partition induces a partition of the array $A$ into square subarrays: for each $Q\in\mathcal{P}(h,w)$, let
$A[Q]\triangleq\bigl(A(i,j)\bigr)_{(i,j)\in Q}.$
With a slight abuse of notation, we use the same symbol $Q$ to refer to both the index square $Q\subseteq\mathcal{I}(h,w)$ and the associated square block $A[Q]$ whenever no confusion arises.
@@
Lemma 4 (Consecutive $\alpha$-exponents after the reshape).
@@
$S^{(1)}$ $=\operatorname{Span}_{\mathbb{E}}\{\overline{R}(0,qs+u):q\in[0,a_{0}-1],u\in[0,s-1]\},$
$S^{(2)}$ $=\operatorname{Span}_{\mathbb{E}}\{\overline{R}(0,a_{0}s+j):0\leq j\leq b_{1}-1\}.$
@@
$\alpha^{7},\alpha^{8},\alpha^{9}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,0)\}_{t=0}^{4},R(0,5),R(1,5)\},$
$\alpha^{9}\beta^{2}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,2)\}_{t=0}^{4},R(2,5),R(3,5)\},$
$\alpha^{8}\beta,\alpha^{9}\beta,\alpha^{10}\beta$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,1)\}_{t=0}^{4},R(0,6),R(1,6)\},$
$\alpha^{10}\beta^{3}$ $\in\operatorname{Span}_{\mathbb{E}}\{\{\overline{R}(t,3)\}_{t=0}^{4},R(2,6),R(3,6)\},$
@@ Appendix A Proof of Proposition 1
Recall that determining $\mathcal{R}_{j}$ amounts to identifying the $\beta$-exponents appearing in the $(a_{0}s+j)$-th column of $R$, where $j\in\{0,1,\dots,b_{1}-1\}$. For simplicity, let $R_{0}$ denote the last $b_{1}$ columns of $R$. The following useful fact follows directly from the definition of the map $\mathsf{Reshape}_{p,s}(\cdot)$.
Fact: After the $\mathsf{Reshape}_{p,s}(\cdot)$, the squares in $R$ corresponding to Euclidean squares in $B$ form an Euclidean square partition of $R$, so does the squares in $R_{0}$, hence we can denote the set of all squares in $R_{0}$ by $\mathcal{P}(s,b_{1})$.
From the proof of Lemma 4, the exponents of $\beta$ inside each column of any Euclidean square $P$ in $R_{0}$ are the same. This motivates the following definition..
@@
Fix $j\in\{0,1,\dots,b_{1}-1\}$. Define $\mathcal{YV}(j)$ to be the set of all pairs $(y,v)$ for which there exist a Euclidean square $Q\in\mathcal{P}(s,b_{1})$ with upper-left corner $(y,x)$ and side length $t$ such that
@@
Equivalently, $(y,v)\in\mathcal{YV}(j)$ if and only if the column indexed by $j$ in $\mathcal{I}(s,b_{1})$ intersects a Euclidean square whose upper-left row coordinate is $y$ with within-square column index $v$.
From Lemma 4, it follows immediately that
@@
Hence the core task is to characterize $\mathcal{YV}(j)$ for each $j$. This is purely a question about the Euclidean square partition $\mathcal{P}(s,b_{1})$. Set $b_{0}:=s$, the recursive construction of $\mathcal{P}(s,b_{1})$ begins with $b_{0}=a_{1}b_{1}+b_{2}$ and alternates between two types of decomposition steps:
@@
    which places $a_{2t-1}$ squares of side length $b_{2t-1}$ along the vertical direction;
@@
    which places $a_{2t}$ squares of side length $b_{2t}$ along the horizontal direction.
These steps are applied to the successive residual rectangles produced in the recursion.
Fix $j\in\{0,1,\dots,b_{1}-1\}$ and track the fixed global column $a_{0}s+j$ through the Euclidean recursion of $\mathcal{P}(s,b_{1})$. At any stage of the recursion, we describe the position of this fixed global column by its local column coordinate within the current residual rectangle. Vertical decomposition steps only cut the residual rectangle along the vertical direction and therefore preserve its horizontal interval, hence they do not change the local column coordinate. In contrast, at the $(2t)$-th horizontal decomposition step induced by $b_{2t-1}=a_{2t}b_{2t}+b_{2t+1}$, the current residual rectangle of width $b_{2t-1}$ is partitioned, from left to right, into $a_{2t}$ squares of side length $b_{2t}$ and one residual sub-rectangle of width $b_{2t+1}$. Accordingly, we define the local column coordinate recursively by
$j^{(0)}:=j,\qquad j^{(2t)}:=j^{(2t-2)}-\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor b_{2t}\quad(t\geq 1),$ (13)
so that $j^{(2t)}$ is the remainder of $j^{(2t-2)}$ upon division by $b_{2t}$.
Key observation: the column $a_{0}s+j$ enters the (2t1)(2t\!-\!1)-st vertical decomposition step if and only if, at each preceding horizontal decomposition step $(2k)$ with $k\in\{1,\dots,t-1\}$, the column lies in the residual sub-rectangle (rather than in one of the $a_{2k}$ squares of side length $b_{2k}$), i.e.,
@@
Then we have:
  • (Vertical case: squares produced at the (2t1)(2t\!-\!1)-st vertical step). At the (2t1)(2t\!-\!1)-st vertical step, the recursion places $a_{2t-1}$ squares of side length $b_{2t-1}$ stacked vertically. Since these squares cover the full current width $b_{2t-1}$, the column intersects all of them, with the same within-square column index $v=j^{(2t-2)}$. Their upper-left row coordinates are
    $y=\Big(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\Big)+q\,b_{2t-1},\qquad q=0,1,\dots,a_{2t-1}-1.$
    Hence $\mathcal{YV}(j)$ contains the $a_{2t-1}$ pairs
    $(y,v)=\Bigg(\Big(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\Big)+q\,b_{2t-1},\ j^{(2t-2)}\Bigg),\qquad q=0,1,\dots,a_{2t-1}-1.$
  • (Horizontal terminating case: the unique square produced at the $(2t)$-th horizontal step). Now consider the subsequent $(2t)$-th horizontal step. If
    $\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor<a_{2t},$
    then the column lies in one of the $a_{2t}$ squares of side length $b_{2t}$ produced in this step, hence intersects exactly one such square. Its upper-left row coordinate equals
    $y=\sum_{k=1}^{t}a_{2k-1}b_{2k-1},$
    and the within-square column index is $v=j^{(2t)}$. Thus $\mathcal{YV}(j)$ contains the additional (terminating) pair
@@
    If instead $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor=a_{2t}$, then the column lies in the residual sub-rectangle and this horizontal step contributes no pair, as the recursion continues to $(2t+1)$-th vertical decomposition.
Collecting the contributions from the vertical case (for all reachable depths $t$) and from the horizontal terminating case (whenever termination occurs at step $(2t)$) yields the following closed form:
$\hskip 73.97733pt+\,\sum_{\begin{subarray}{c}1\leq t\leq\lfloor m/2\rfloor\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor\frac{j^{(2k-2)}}{b_{2k}}\right\rfloor=a_{2k}\\ \left\lfloor\frac{j^{(2t-2)}}{b_{2t}}\right\rfloor<a_{2t}\end{subarray}}\beta^{\left(\sum_{k=1}^{t}a_{2k-1}b_{2k-1}\right)+j^{(2t)}}\Bigg),$
which finished the proof of Proposition 1.

### v2

@@ II-C Reed-Solomon codes as MSR codes
Let $E$ be an extension field of a finite field $F$ with $E:F]=\ell$. Assume $\{\zeta_{1},\dots,\zeta_{\ell}\}$ and $\{\mu_{1},\dots,\mu_{\ell}\}$ are dual bases of $E$ over $F$, i.e., ${\rm tr}(\zeta_{i}\mu_{j})=\delta_{ij}$, where ${\rm tr}(\cdot)$ denotes the trace function from $E$ to $F$ and $\delta_{ij}$ is the Kronecker delta. Then for any $\gamma\in E$,
@@
For a vector space $V$ over a field $F$ and a set of vectors $A=\left(a_{1},\ldots,a_{\ell}\right)\subset V$, define
$\operatorname{Span}_{F}(A)=\left\{\sum_{i=1}^{\ell}\gamma_{i}a_{i}:\ \gamma_{i}\in F\right\}.$
@@
    Let $h(x)=\prod_{j\in[n]\setminus(R_{i}\cup\{i\})}(x-\alpha_{j})$, then for $0\leq t\leq s-1$, $(v_{1}\alpha_{1}^{t}h(\alpha_{1}),\dots,v_{n}\alpha_{n}^{t}h(\alpha_{n}))\in\mathcal{C}^{\perp},$ hence $\sum_{j=1}^{n}c_{j}v_{j}\alpha_{j}^{t}h(\alpha_{j})=0$
@@
    Assume that $R_{i}=\{i_{1},\dots,i_{d}\}$. Then for each $0\leq t\leq s-1$,
    $c_{i}v_{i}\alpha_{i}^{t}h(\alpha_{i})+\sum_{j=1}^{d}c_{i_{j}}v_{i_{j}}\alpha_{i_{j}}^{t}h(\alpha_{i_{j}})=0.$
@@
    can be used to recover $c_{i}$, as $\left\{e_{m}v_{i}\alpha_{i}^{t}h(\alpha_{i})\right\}_{0\leq t\leq s-1,1\leq m\leq p_{i}}$ forms a basis of $\mathbb{E}$ over $F_{i}$.
@@
    Hence, from each helper node (for example, the $i_{1}$-th node in $R_{i}$), one only needs to compute $\{{\rm tr}_{i}(c_{i_{1}}v_{i_{1}}e_{m})\}_{1\leq m\leq p_{i}}$, i.e., $p_{i}$ symbols from $F_{i}$. Consequently, the total repair bandwidth is
    $dp_{i}=\frac{sp_{i}d}{d+1-k},$
    which meets the cut-set bound (1).
@@ III Main results
In this section, we present our main results, which show that the congruence condition $p_{i}\equiv 1\pmod{s}$ in Theorem 2 can be removed. As a result, we obtain an RS–MSR code with smaller subpacketization $\ell=s\prod_{i=1}^{n}p_{i}$ for any choice of distinct primes $p_{1},\ldots,p_{n}>s$.
From the discussion in Section II, the key is to show the existence of an $F_{i}$-subspace $S_{i}\subseteq\mathbb{E}$ that satisfies (3) without requiring the condition $p_{i}\equiv 1\pmod{s}$. Equivalently,
Lemma 3 (Existence of an $\mathbb{F}$-subspace with $\alpha$-layer decomposition).
@@
Figure 1: Original $p\times s$ array
@@
Figure 2: Target $s\times p$ array
@@
where $b_{m}=\gcd(p,s)$. We define a deterministic partition of any array of size $p\times s$ by partitioning its index set $\mathcal{I}(p,s)\triangleq[0,p-1]\times[0,s-1]$ into disjoint square blocks via the following recursive procedure.
@@
Let $h,w$ be positive integers and consider the rectangle $\mathcal{I}(h,w)=[0,h-1]\times[0,w-1]$. Let $t:=\min\{h,w\}$. Tile $\mathcal{I}(h,w)$ by placing as many $t\times t$ squares as possible along the longer side (left-to-right if $w\geq h$, and top-to-bottom if $h>w$). This leaves a (possibly empty) residual rectangle $\mathcal{I}(h^{\prime},w^{\prime})$; if it is nonempty, apply the same procedure recursively to $\mathcal{I}(h^{\prime},w^{\prime})$. The resulting collection of squares forms a disjoint partition of $\mathcal{I}(h,w)$, called the Euclidean square partition of $\mathcal{I}(h,w)$, and is denoted as $\mathcal{P}(h,w)$. Each square in $\mathcal{P}(h,w)$ is called a Euclidean square.
@@
Lemma 4 (Consecutive $\alpha$-exponents in each reshaped column).
@@
$S^{(1)}$ $=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(0,qs+u):q\in[0,a_{0}-1],u\in[0,s-1]\},$
$S^{(2)}$ $=\operatorname{Span}_{\mathbb{F}}\{\overline{R}(0,a_{0}s+j):0\leq j\leq b_{1}-1\}.$
@@
$\alpha^{7},\alpha^{8},\alpha^{9}$ $\in\operatorname{Span}_{\mathbb{F}}\{\{\overline{R}(t,0)\}_{t=0}^{4},R(0,5),R(1,5)\},$
$\alpha^{9}\beta^{2}$ $\in\operatorname{Span}_{\mathbb{F}}\{\{\overline{R}(t,2)\}_{t=0}^{4},R(2,5),R(3,5)\},$
$\alpha^{8}\beta,\alpha^{9}\beta,\alpha^{10}\beta$ $\in\operatorname{Span}_{\mathbb{F}}\{\{\overline{R}(t,1)\}_{t=0}^{4},R(0,6),R(1,6)\},$
$\alpha^{10}\beta^{3}$ $\in\operatorname{Span}_{\mathbb{F}}\{\{\overline{R}(t,3)\}_{t=0}^{4},R(2,6),R(3,6)\}.$
@@ Appendix A Proof of Proposition 1
Recall that determining $\mathcal{R}_{j}$ amounts to identifying all $\beta$-exponents appearing in column $a_{0}s+j$ of the reshaped array $R$, where $j\in\{0,1,\dots,b_{1}-1\}$. Denote by $R_{\mathrm{1}}$ the subarray of $R$ consisting of its last $b_{1}$ columns. By Definition 2, the map $\mathsf{Reshape}_{p,s}(\cdot)$ preserves the Euclidean square partition structure; consequently, the squares in $R_{\mathrm{1}}$ induced by $\mathcal{P}(p,s)$ form precisely the Euclidean square partition $\mathcal{P}(s,b_{1})$ of the index set $\mathcal{I}(s,b_{1}):=[0,s-1]\times[0,b_{1}-1]$.
From Lemma 4, the $\beta$-exponent is constant along each column of any Euclidean square in $R_{\mathrm{1}}$. This structural property leads to the following characterization.
@@
Fix $j\in\{0,1,\dots,b_{1}-1\}$. Define $\mathcal{YV}(j)$ as the set of pairs $(y,v)$ for which there exists a Euclidean square $Q\in\mathcal{P}(s,b_{1})$ with top-left corner $(y,x)$ and side length $t$ satisfying
@@
Equivalently, $(y,v)\in\mathcal{YV}(j)$ if and only if column $j$ of $\mathcal{I}(s,b_{1})$ intersects a square whose top-left row coordinate is $y$ at local column index $v$ within that square.
By Lemma 4, we obtain the following reduction.
@@
Thus the problem reduces to a combinatorial characterization of $\mathcal{YV}(j)$. Let $b_{0}:=s$ and consider the Euclidean algorithm applied to $(b_{0},b_{1})$:
$b_{0}=a_{1}b_{1}+b_{2},\quad b_{1}=a_{2}b_{2}+b_{3},\quad\dots,\quad b_{m-2}=a_{m-1}b_{m-1}+b_{m},\quad b_{m-1}=a_{m}b_{m},$ (11)
where $b_{m}=\gcd(s,b_{1})$ and $m\geq 1$ denotes the total number of division steps. The partition $\mathcal{P}(s,b_{1})$ is constructed recursively by alternating between two types of decomposition:
@@
    which tiles the current residual rectangle of height $b_{2t-2}$ with $a_{2t-1}$ squares of side length $b_{2t-1}$ stacked vertically, leaving a residual rectangle of height $b_{2t}$.
@@
    which tiles the current residual rectangle of width $b_{2t-1}$ with $a_{2t}$ squares of side length $b_{2t}$ placed horizontally, leaving a residual rectangle of width $b_{2t+1}$.
Fix $j\in\{0,1,\dots,b_{1}-1\}$ and track column $a_{0}s+j$ through this recursion. At stage $2t$ of the recursion, let $j^{(2t)}$ denote the column index of $a_{0}s+j$ within the current residual rectangle. Since vertical steps preserve the horizontal extent of the residual rectangle, $j^{(2t)}$ changes only at horizontal steps. Specifically, we define recursively
$j^{(0)}:=j,\qquad j^{(2t)}:=j^{(2t-2)}\pmod{b_{2t}}=j^{(2t-2)}-\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor b_{2t},\quad t\geq 1.$ (14)
The recursion proceeds deterministically. Since the Euclidean algorithm consists of exactly $m$ division steps, the trajectory of column $a_{0}s+j$ terminates at the first horizontal decomposition step $(2t)$ for which $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor<a_{2t}$, or exhausts the entire chain after at most $\lfloor m/2\rfloor$ horizontal decomposition steps. In either case, termination occurs after finitely many steps. Every intersection between column $a_{0}s+j$ and a Euclidean square in $\mathcal{P}(s,b_{1})$ occurs precisely at one of the following mutually exclusive events:
  1. (i)
    The column reaches a vertical decomposition step (2t1)(2t\!-\!1), i.e., it remains in the residual sub-rectangle after each preceding horizontal step $(2k)$ for $k=1,\dots,t-1$:
@@
    At such a step, the column intersects all $a_{2t-1}$ vertically stacked squares (since they cover the full width $b_{2t-1}$ of the current residual rectangle), yielding $a_{2t-1}$ distinct pairs in $\mathcal{YV}(j)$.
  2. (ii)
    The column terminates at a horizontal decomposition step $(2t)$, i.e., condition (15) holds for $k=1,\dots,t-1$ but
    $\Big\lfloor\frac{j^{(2t-2)}}{b_{2t}}\Big\rfloor<a_{2t}.$
    In this case the column lies within one of the $a_{2t}$ complete squares (rather than the residual sub-rectangle) and intersects exactly one square at this step, yielding a single additional pair in $\mathcal{YV}(j)$.
No other intersections occur: when $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor=a_{2t}$ at a horizontal step, the column lies in the residual sub-rectangle of width $b_{2t+1}$ and does not intersect any square produced at that step; such intersections are accounted for in subsequent recursion steps. Consequently, $\mathcal{YV}(j)$ is precisely the disjoint union of contributions from events (i) over all $t$ satisfying (15) and from event (ii) at the unique terminating horizontal step (if any).
Under condition (15), the contributions to $\mathcal{YV}(j)$ are as follows:
  • Vertical contributions (step 2t12t\!-\!1). The column intersects all $a_{2t-1}$ vertically stacked squares of side length $b_{2t-1}$. Each such square has top-left row coordinate
    $y=\Big(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\Big)+q\,b_{2t-1},\quad q=0,1,\dots,a_{2t-1}-1,$
    and local column index $v=j^{(2t-2)}$ within the square. Hence $\mathcal{YV}(j)$ contains the pairs
    $(y,v)=\Bigg(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}+q\,b_{2t-1},\ j^{(2t-2)}\Bigg),\quad q=0,\dots,a_{2t-1}-1.$
  • Horizontal terminating contribution (step $2t$). If condition (15) holds for $k=1,\dots,t-1$ but $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor<a_{2t}$, then the column lies within one of the $a_{2t}$ squares of side length $b_{2t}$ and contributes exactly one pair with top-left row coordinate
    $y=\sum_{k=1}^{t}a_{2k-1}b_{2k-1}$
    and local column index $v=j^{(2t)}$ within that square, namely
@@
    If instead $\big\lfloor j^{(2t-2)}/b_{2t}\big\rfloor=a_{2t}$, the column remains in the residual rectangle; no square is intersected at this horizontal step, and the recursion proceeds to step $(2t+1)$.
The closed-form expression for $\mathcal{R}_{j}$ aggregates precisely these two types of contributions:
  • The first summation collects contributions from all vertical decomposition steps (2t1)(2t\!-\!1) that the column reaches (i.e., all $t$ satisfying (15)). At each such step, the column intersects all $a_{2t-1}$ squares produced, yielding $a_{2t-1}$ terms.
  • The second summation captures the contribution from the unique terminating horizontal step $(2t)$ (if any), where the column falls into one of the $a_{2t}$ complete squares rather than the residual sub-rectangle, yielding exactly one additional term.
No explicit partition of the range of $j$ is required: the constraints in the summation indices automatically select the relevant steps for each $j$ according to its trajectory through the Euclidean recursion.
Aggregating all vertical contributions for indices $t$ satisfying (15) and the unique horizontal terminating contribution (if any) yields the closed-form expression:
$\mathcal{R}_{j}$ $=\alpha^{a_{0}s+j}\Bigg(\sum_{\begin{subarray}{c}1\leq t\leq\lceil m/2\rceil\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor j^{(2k-2)}/b_{2k}\right\rfloor=a_{2k}\end{subarray}}\ \sum_{q=0}^{a_{2t-1}-1}\beta^{\left(\sum_{k=1}^{t-1}a_{2k-1}b_{2k-1}\right)+q\,b_{2t-1}+j^{(2t-2)}}$
$\hskip 73.97733pt+\,\sum_{\begin{subarray}{c}1\leq t\leq\lfloor m/2\rfloor\\ \forall k\in\{1,\dots,t-1\}:\ \left\lfloor j^{(2k-2)}/b_{2k}\right\rfloor=a_{2k}\\ \left\lfloor j^{(2t-2)}/b_{2t}\right\rfloor<a_{2t}\end{subarray}}\beta^{\left(\sum_{k=1}^{t}a_{2k-1}b_{2k-1}\right)+j^{(2t)}}\Bigg),$ (16)
which completes the proof of Proposition 1.